# LLaMAX: Scaling Linguistic Horizons of LLM by Enhancing Translation Capabilities Beyond 100 Languages

### v1

@@ 4 Benchmarking Results
While LLaMAX2 has achieved the state-of-the-art translation performance among open-source decoder-only LLMs, the next critical question is whether we can close the gap between LLMs and specialized encoder-decoder translation systems. Table 4 provides a comprehensive comparison, reveals LLaMAX2 has reached the level of the M2M-100-12B model. Future work will be needed to optimize the language extension framework to match the performance of advanced translation systems, e.g., MADLAD-400.
@@
| Direct | BLEU | COMET | ||||||
| zhx | 10.14 | 11.34 | 14.17 | 15.54 | 64.65 | 66.61 | 73.54 | 74.74 |
| xzh | 13.17 | 15.37 | 13.53 | 15.11 | 73.56 | 75.66 | 75.52 | 77.21 |
| dex | 13.62 | 14.24 | 18.96 | 19.38 | 64.67 | 65.79 | 73.82 | 74.36 |
| xde | 16.82 | 18.08 | 19.26 | 20.71 | 71.59 | 73.11 | 74.47 | 76.04 |
| arx | 11.27 | 12.60 | 15.82 | 17.10 | 63.45 | 65.33 | 72.00 | 73.17 |
| xar | 10.20 | 10.88 | 15.32 | 16.00 | 66.49 | 69.54 | 75.40 | 76.32 |
| nex | 7.96 | 10.29 | 14.49 | 16.16 | 62.95 | 67.87 | 74.64 | 76.86 |
| xne | 3.83 | 7.08 | 15.47 | 16.86 | 46.56 | 58.89 | 67.36 | 69.47 |
| azx | 6.98 | 9.52 | 11.34 | 13.54 | 60.61 | 65.16 | 70.91 | 73.60 |
| xaz | 4.81 | 6.96 | 10.27 | 11.44 | 58.30 | 67.52 | 72.03 | 75.60 |
| cebx | 8.52 | 10.69 | 15.53 | 16.98 | 55.26 | 60.71 | 68.67 | 70.76 |
| xceb | 4.18 | 7.17 | 16.11 | 18.94 | 52.68 | 59.55 | 65.05 | 66.52 |
| Avg. | 9.29 | 11.19 | 15.02 | 16.48 | 61.73 | 66.31 | 71.95 | 73.72 |
LLaMAX circumvents catastrophic forgetting issue.
A common concern with continual pre-training on additional multilingual corpus is that the process might disturb the parametric knowledge and working pattern of the original model, a phenomenon known as catastrophic forgetting Goodfellow et al. (2013). Furthermore, we compare LLaMAX2 with LLaMA2 on popular English benchmarks that measure a diverse set of core capabilities of LLMs. Experiment results in Table 6 show that the two models achieve very similar performance on these benchmarks (More details about these benchmarks are in Appendix B.), demonstrating that our continual pre-training does not compromise the general capability of the base model.
Comparsion between LLaMAX2-Alpaca and language-specific LLMs.
In addition, we conduct a deeper analysis of translation output to identify the factors contributing to the limited improvement in translation performance. The experimental results in Table 7 indicate that the language-specific LLM obtained through the traditional pipeline tends to output specific languages, while LLaMAX2 can accurately produce the answer with the corresponding language. We provide more analysis on Swallow in Appendix I, and arrive at the same conclusion.
5 Related Work
Multilingual Large Language Models.
Large Language Model (LLMs; OpenAI, 2023; Zhang et al., 2022; Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023a, b) trained with English-centric data can also solve various non-English tasks (Hendrycks et al., 2021a, b; Srivastava et al., 2022; Kwiatkowski et al., 2019; Hendrycks et al., 2021c), but the performance between non-English and English is significantly large Yuan et al. (2023b). Efforts to develop more multilingual LLMs in two different ways: retraining LLMs with diverse multilingual data from scratch (Wei et al., 2023); or continuous training of pre-trained models using language-specific data with the option to expand the vocabulary (Zhao et al., 2024a; Cui et al., 2024; Faysse et al., 2024; Alves et al., 2024). Instead of training from scratch, continual pre-training aims at updating pre-trained models with new data, making the process more efficient and cost-effective (Gupta et al., 2023; Alves et al., 2024; Xie et al., 2023).
Multilinguality in LLMs.
Recent research has shed light on the multilingual capabilities of LLMs. A comprehensive survey by Huang et al. (2024) discusses various aspects of multilingualism in LLMs, including training and inference methods, model security, multi-domain with languages culture, and emphasizes the need for language-fai technology. Yuan et al. (2023b) analysis multilingualism of LLMs from the vocabulary sharing aspect. Zhao et al. (2024b) delve into the architecture of LLMs to find how LLMs handle multilingualism. Recently, Li et al. (2024) quantify the multilingual performance of LLMs. These studies provide valuable insights into the multilingual capabilities of LLMs, and the key technical design of continual pre-training for LLaMAX.
6 Conclusion
Acknowledgments
Authors of this paper would like to thank Zixian Huang, Qiushi Sun, Fangzhi Xu, Hanxu Hu, Chuanyang Jin, Yichao Du, and Zichen Ding for giving many helpful comments on previous versions of this paper.
Appendix A Limitations
This work focuses on the discussion of some key technologies, such as the use of vocabulary lists and the determination of data augmentation schemes. However, it does not delve into further processing of the quality of open-source data. We acknowledge a gap in the literature regarding the thorough evaluation of open-source data quality, suggesting an opportunity for future research to improve data preprocessing methods for better model training outcomes.
Appendix B Data Information
| om | Oromo | 18,895 | 191,319 | 96 | or | Odia | 100,530 | 812,235 | 97 | ||
| sl | Slovenian | 731,613 | 4,073,213 | 97 | umb | Umbundu | 0 | 54 | 2 | ||
| is | Icelandic | 704,159 | 4,088,886 | 97 | kam | Kamba | 0 | 8 | 1 | ||
| sv | Swedish | 726,893 | 4,213,939 | 97 | ff | Fulani | 26 | 313,870 | 97 | ||
In this section, we will introduce the sources of our training data (Section B.1), the evaluation benchmarks (Section B.2). For translation tasks, we apply beam search to each model with beam size=4.
B.1 Training Dataset
The dataset was compiled from three distinct open-source datasets, with details on data statistics and supported languages presented in the Table 9.
MC4 (Xue et al., 2021)
MADLAD-400 (Kudugunta et al., 2024)
Lego-MT Yuan et al. (2023a)
B.2 Evaluation Benchmark
Flores-101 (Goyal et al., 2022)
Flores-200 (Team et al., 2022)
WMT-23 (Kocmi and Federmann, 2023)
TICO Anastasopoulos et al. (2020)
TED Cettolo et al. (2012)
X-CSQA Lin et al. (2021a)
XStoryCloze (Lin et al., 2021b)
| enx | xen | tax | xta | thx | xth | zhx | xzh | ||
| ++ | MUSE: 1-hop | 18.80 | 26.56 | 4.78 | 1.79 | 7.31 | 3.18 | 11.35 | 7.28 |
| ++ | MUSE: 2-hop | 18.70 | 26.50 | 4.47 | 1.83 | 7.08 | 3.26 | 10.74 | 6.68 |
| ++ | PanLex: 1-hop | 19.33 | 26.54 | 4.40 | 1.83 | 7.57 | 3.31 | 10.86 | 8.08 |
| 1-hop translation | 2-hop translation | ||
| Direction | Example | Direction | Example |
| enfr | dog chien | enfrde | dog chien Hund |
| frde | chien Hund | ||
XCOPA (Ponti et al., 2020)
XWinograd (Muennighoff et al., 2022; Tikhonov and Ryabinin, 2021)
XNLI (Conneau et al., 2018)
MGSM (Shi et al., 2023)
MMLU (Hendrycks et al., 2021a, b)
BBH (Srivastava et al., 2022)
HellaSwag (Zellers et al., 2019)
WinoG (Sakaguchi et al., 2021)
NQ (Kwiatkowski et al., 2019)
HumanEval (Chen et al., 2021)
MBPP (Austin et al., 2021)
GSM8K (Cobbe et al., 2021)
Math (Hendrycks et al., 2021c)
Appendix C Model Information
| cebx | xceb | QNLI | QQP | MRPC | XStoryCloze | XCOPA | XWinograd | |
| Setting | cebca | cebde | ceben | cebes | cebfr | cebit | cebpt | cebru |
| Setting | caceb | deceb | enceb | esceb | frceb | itceb | ptceb | ruceb |
Model details about the baseline models for comparison, including decode-only large language models (LLMs) in Section C.1 as well as translation models in Section C.2 with an encoder-decoder structure.
C.1 Large Language Models
LLaMA2 (Touvron et al., 2023b)
LLaMAX2
PolyLM (Wei et al., 2023)
Yayi2 (Luo et al., 2023)
TowerInstruct (Alves et al., 2024)
Aya-23 (Aryabumi et al., 2024)
ChineseLLaMA2-Alpaca Cui et al. (2024)
LLaMA2-SFT (Taori et al., 2023)
Qwen2-7B-Instruct (Bai et al., 2023)
Swallow (Fujii et al., 2024)
C.2 Translation Models
@@
Appendix D The correlation between fertility and representation quality.
@@
Appendix E Introduction to KS-Lottery.
@@
Appendix F 1-hop translation in data augmentation is enough.
@@
Appendix G Design of parallel format
@@
Appendix H Comparison Results Between Our Model and GPT-4
@@
Appendix I Comparison LLaMAX2-Alpaca with language-specific LLMs.
@@
Appendix J Prompt Templates
We offer a comprehensive collection of prompt instruction templates, as illustrated in Table D, which are utilized for all evaluated LLMs. These templates are meticulously designed based on existing LLMs, playing a crucial role in obtaining accurate model results and ensuring fairness in comparisons. Our goal in providing these templates is to promote transparency and make it easier to reproduce our findings.

### v2

@@ 4 Benchmarking Results
While LLaMAX2 has achieved the state-of-the-art translation performance among open-source decoder-only LLMs, the next critical question is whether we can close the gap between LLMs and specialized encoder-decoder translation systems. Table 3 provides a comprehensive comparison, reveals LLaMAX2 has reached the level of the M2M-100-12B model. Future work will be needed to optimize the language extension framework to match the performance of advanced translation systems.
| Direct | spBLEU | COMET | ||||||
| zhX | 10.14 | 11.34 | 14.17 | 15.54 | 64.65 | 66.61 | 73.54 | 74.74 |
| Xzh | 13.17 | 15.37 | 13.53 | 15.11 | 73.56 | 75.66 | 75.52 | 77.21 |
| deX | 13.62 | 14.24 | 18.96 | 19.38 | 64.67 | 65.79 | 73.82 | 74.36 |
| Xde | 16.82 | 18.08 | 19.26 | 20.71 | 71.59 | 73.11 | 74.47 | 76.04 |
| arX | 11.27 | 12.60 | 15.82 | 17.10 | 63.45 | 65.33 | 72.00 | 73.17 |
| Xar | 10.20 | 10.88 | 15.32 | 16.00 | 66.49 | 69.54 | 75.40 | 76.32 |
| neX | 7.96 | 10.29 | 14.49 | 16.16 | 62.95 | 67.87 | 74.64 | 76.86 |
| Xne | 3.83 | 7.08 | 15.47 | 16.86 | 46.56 | 58.89 | 67.36 | 69.47 |
| azX | 6.98 | 9.52 | 11.34 | 13.54 | 60.61 | 65.16 | 70.91 | 73.60 |
| Xaz | 4.81 | 6.96 | 10.27 | 11.44 | 58.30 | 67.52 | 72.03 | 75.60 |
| cebX | 8.52 | 10.69 | 15.53 | 16.98 | 55.26 | 60.71 | 68.67 | 70.76 |
| Xceb | 4.18 | 7.17 | 16.11 | 18.94 | 52.68 | 59.55 | 65.05 | 66.52 |
| Avg. | 9.29 | 11.19 | 15.02 | 16.48 | 61.73 | 66.31 | 71.95 | 73.72 |
LLaMAX circumvents catastrophic forgetting issue.
A common concern with continual pre-training on additional multilingual corpus is that the process might disturb the parametric knowledge and working pattern of the original model, a phenomenon known as catastrophic forgetting Goodfellow et al. (2013). Furthermore, we compare LLaMAX2 with LLaMA2 on popular English benchmarks that measure a diverse set of core capabilities of LLMs. Experiment results in Table 5 show that the two models achieve very similar performance on these benchmarks (More details about these benchmarks are in Appendix A.), demonstrating that our continual pre-training does not compromise the general capability of the base model.
5 Related Work
Multilingual Large Language Models. Large Language Model (LLMs; OpenAI, 2023; Zhang et al., 2022; Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023a, b) trained with English-centric data can also solve various non-English tasks (Hendrycks et al., 2021a, b; Srivastava et al., 2022; Kwiatkowski et al., 2019; Hendrycks et al., 2021c), but the performance between non-English and English is significantly large Yuan et al. (2024b). Efforts to develop more multilingual LLMs in two different ways: retraining LLMs with diverse multilingual data from scratch (Wei et al., 2023); or continuous training of pre-trained models using language-specific data with the option to expand the vocabulary (Zhao et al., 2024a; Cui et al., 2024; Faysse et al., 2024; Alves et al., 2024). Instead of training from scratch, continual pre-training aims at updating pre-trained models with new data, making the process more efficient and cost-effective (Gupta et al., 2023; Alves et al., 2024; Xie et al., 2023).
Multilinguality in LLMs. Recent research has shed light on the multilingual capabilities of LLMs. A comprehensive survey by Huang et al. (2024a) discusses various aspects of multilingualism in LLMs, including training and inference methods, model security, multi-domain with languages culture, and emphasizes the need for language-fai technology. Yuan et al. (2024b) analysis multilingualism of LLMs from the vocabulary sharing aspect. Zhao et al. (2024b) delve into the architecture of LLMs to find how LLMs handle multilingualism. Recently, Li et al. (2024) quantify the multilingual performance of LLMs. These studies provide valuable insights into the multilingual capabilities of LLMs, and the key technical design of continual pre-training for LLaMAX.
6 Conclusion
Acknowledgments
Authors of this paper would like to thank Zixian Huang, Qiushi Sun, Fangzhi Xu, Hanxu Hu, Chuanyang Jin, Yichao Du, and Zichen Ding for giving many helpful comments on previous versions of this paper. We deeply express our gratitude to Shanghai AI Laboratory. This work is partially supported by the National Key R&D Program of China (NO.2022ZD0160100).
Outline
Section D: A detailed introduction to the KS-Lottery method.
Appendix A Data Information
| om | Oromo | 18,895 | 191,319 | 96 | or | Odia | 100,530 | 812,235 | 97 | ||
| sl | Slovenian | 731,613 | 4,073,213 | 97 | umb | Umbundu | 0 | 54 | 2 | ||
| is | Icelandic | 704,159 | 4,088,886 | 97 | kam | Kamba | 0 | 8 | 1 | ||
| sv | Swedish | 726,893 | 4,213,939 | 97 | ff | Fulani | 26 | 313,870 | 97 | ||
| ISO | # Para. | # Mono. | ISO | # Para. | # Mono. | ISO | # Para. | # Mono. | ISO | # Para. | # Mono. |
| af | 201,367,199 | 360,215,552 | hi | 593,592,809 | 366,433,792 | mn | 332,967,182 | 359,067,648 | tg | 347,063,556 | 358,521,344 |
| am | 903,470,803 | 137,815,552 | hr | 212,708,153 | 153,600,000 | mr | 609,131,394 | 359,619,584 | th | 597,728,017 | 362,352,128 |
| ar | 1,054,714,212 | 366,624,256 | hu | 232,631,392 | 374,517,248 | ms | 234,113,298 | 364,490,240 | tl | 244,687,985 | 0 |
| as | 313,146,729 | 17,318,400 | hy | 579,250,350 | 364,971,520 | mt | 102,804,684 | 343,918,592 | tr | 272,252,613 | 372,588,032 |
| ast | 70,949,987 | 0 | id | 232,953,602 | 362,476,544 | my | 1,002,285,410 | 296,529,920 | uk | 218,572,425 | 365,660,672 |
| az | 654,492,231 | 365,028,864 | ig | 242,306,836 | 75,427,328 | ne | 1,237,255,500 | 359,595,008 | umb | 3,170 | 0 |
| be | 306,891,318 | 362,643,456 | is | 251,875,378 | 360,529,408 | nl | 193,189,558 | 394,126,336 | ur | 557,337,279 | 364,213,248 |
| bg | 229,686,547 | 364,288,000 | it | 195,146,279 | 433,206,784 | no | 190,141,837 | 369,518,080 | uz | 148,867,813 | 370,225,152 |
| bn | 755,297,957 | 362,034,688 | ja | 292,857,869 | 371,944,960 | ns | 6,056,515 | 0 | vi | 372,555,263 | 389,361,664 |
| bs | 155,671,162 | 153,600,000 | jv | 150,347,166 | 258,876,928 | ny | 194,642,682 | 116,193,280 | wo | 45,422,689 | 445,952 |
| ca | 196,058,689 | 370,993,664 | ka | 627,397,650 | 360,199,680 | oc | 91,504,042 | 18,626,048 | xh | 242,467,614 | 62,832,640 |
| ceb | 135,958,864 | 214,045,696 | kam | 477 | 0 | om | 13,239,275 | 9,674,240 | yo | 282,242,956 | 50,319,872 |
| cs | 218,791,438 | 364,123,648 | kea | 0 | 0 | or | 289,074,437 | 51,471,360 | zh | 878,523,029 | 371,769,344 |
| cy | 247,455,922 | 360,195,584 | kk | 299,995,454 | 359,346,688 | pa | 1,088,815,314 | 263,161,344 | zhtrad | 252,942,548 | 0 |
| da | 201,340,176 | 369,430,016 | km | 1,266,785,006 | 352,097,280 | pl | 223,440,053 | 405,928,448 | zu | 189,932,839 | 240,846,336 |
| de | 456,147,707 | 451,355,136 | kn | 1,198,503,370 | 364,969,984 | ps | 482,900,652 | 301,230,080 | |||
| el | 629,383,799 | 362,368,512 | ko | 415,419,459 | 364,239,872 | pt | 189,507,878 | 407,458,816 | |||
| en | 523,902,024 | 433,516,544 | ku | 494,346,376 | 264,826,368 | ro | 224,472,825 | 359,425,024 | |||
| es | 193,760,704 | 361,629,184 | ky | 284,261,983 | 360,672,256 | ru | 213,581,742 | 436,944,384 | |||
| et | 223,902,240 | 361,840,640 | lb | 58,408,912 | 293,972,992 | sd | 107,023,879 | 241,775,104 | |||
| fa | 505,307,774 | 369,309,184 | lg | 12,860,033 | 6,671,360 | sk | 232,485,422 | 366,356,480 | |||
| ff | 16,917,593 | 13,312 | ln | 8,942,304 | 1,702,400 | sl | 211,807,076 | 374,585,856 | |||
| fi | 242,982,346 | 368,340,992 | lo | 932,379,487 | 183,172,096 | sn | 196,567,944 | 197,933,056 | |||
| fr | 198,627,139 | 403,105,792 | lt | 231,673,345 | 367,811,584 | so | 255,665,827 | 357,306,368 | |||
| ga | 190,006,560 | 355,051,520 | luo | 4,996 | 0 | sr | 217,789,020 | 364,305,920 | |||
| gl | 145,312,272 | 371,974,144 | lv | 261,558,146 | 358,855,168 | sv | 190,891,437 | 372,169,216 | |||
| gu | 1,157,006,948 | 360,764,928 | mi | 234,795,047 | 92,507,136 | sw | 218,972,852 | 358,616,064 | |||
| ha | 185,399,766 | 215,533,568 | mk | 230,161,774 | 359,441,920 | ta | 805,830,274 | 364,473,856 | |||
| he | 401,537,464 | 153,600,000 | ml | 773,141,254 | 366,278,144 | te | 1,387,859,988 | 362,731,008 |
In this section, we will introduce the sources of our training data (Section A.1), the evaluation benchmarks (Section A.2). For translation tasks, we apply beam search to each model with beam size=4.
A.1 Training Dataset
The dataset was compiled from three distinct open-source datasets, with details on supported languages presented in Table 7 and continual pre-training data statistics in Table 7 and Table 8.
MC4 (Xue et al., 2021)
MADLAD-400 (Kudugunta et al., 2024)
Lego-MT Yuan et al. (2023)
A.2 Evaluation Benchmark
Flores-101 (Goyal et al., 2022)
Flores-200 (Team et al., 2022)
WMT-23 (Kocmi and Federmann, 2023)
TICO Anastasopoulos et al. (2020)
TED Cettolo et al. (2012)
X-CSQA Lin et al. (2021a)
XStoryCloze (Lin et al., 2021b)
XCOPA (Ponti et al., 2020)
XWinograd (Muennighoff et al., 2022; Tikhonov and Ryabinin, 2021)
XNLI (Conneau et al., 2018)
MGSM (Shi et al., 2023)
MMLU (Hendrycks et al., 2021a, b)
BBH (Srivastava et al., 2022)
HellaSwag (Zellers et al., 2019)
WinoG (Sakaguchi et al., 2021)
NQ (Kwiatkowski et al., 2019)
HumanEval (Chen et al., 2021)
MBPP (Austin et al., 2021)
GSM8K (Cobbe et al., 2021)
| enX | Xen | taX | Xta | thX | Xth | zhX | Xzh | ||
| LLaMA2 | ✗ | 18.31 | 23.61 | 0.99 | 0.49 | 4.83 | 1.15 | 10.02 | 7.35 |
| ✗ | 19.06 | 25.98 | 3.20 | 0.91 | 7.66 | 3.13 | 11.32 | 7.83 | |
| + | ✗ | 19.46 | 26.40 | 4.17 | 1.76 | 7.28 | 3.02 | 11.65 | 8.82 |
| + | ✗ | 19.22 | 25.91 | 3.51 | 1.34 | 7.64 | 2.83 | 11.56 | 7.99 |
| ++ | ✗ | 19.36 | 26.47 | 4.35 | 1.82 | 7.78 | 3.49 | 11.44 | 9.14 |
| + | ✓ | 19.47 | 26.65 | 4.54 | 1.83 | 7.66 | 3.13 | 11.89 | 9.17 |
| + | ✓ | 18.59 | 25.98 | 3.61 | 1.36 | 6.72 | 2.35 | 10.81 | 6.45 |
| ++ | ✓ | 19.70 | 26.71 | 4.68 | 1.82 | 8.21 | 3.65 | 12.05 | 9.28 |
| ++ | ✓ | 19.17 | 26.58 | 4.57 | 1.95 | 7.12 | 3.12 | 11.52 | 7.73 |
| ++ | ✓ | 18.80 | 26.56 | 4.78 | 1.79 | 7.31 | 3.18 | 11.35 | 7.28 |
| enx | xen | tax | xta | thx | xth | zhx | xzh | ||
| ++ | MUSE: 1-hop | 18.80 | 26.56 | 4.78 | 1.79 | 7.31 | 3.18 | 11.35 | 7.28 |
| ++ | MUSE: 2-hop | 18.70 | 26.50 | 4.47 | 1.83 | 7.08 | 3.26 | 10.74 | 6.68 |
| ++ | PanLex: 1-hop | 19.33 | 26.54 | 4.40 | 1.83 | 7.57 | 3.31 | 10.86 | 8.08 |
| 1-hop translation | 2-hop translation | ||
| Direction | Example | Direction | Example |
| enfr | dog chien | enfrde | dog chien Hund |
| frde | chien Hund | ||
Math (Hendrycks et al., 2021c)
Appendix B Model Information
| cebx | xceb | QNLI | QQP | MRPC | XStoryCloze | XCOPA | XWinograd | |
| Setting | cebca | cebde | ceben | cebes | cebfr | cebit | cebpt | cebru |
| Setting | caceb | deceb | enceb | esceb | frceb | itceb | ptceb | ruceb |
Model details about the baseline models for comparison, including decode-only large language models (LLMs) in Section B.1 as well as translation models in Section B.2 with an encoder-decoder structure.
B.1 Large Language Models
LLaMA2 (Touvron et al., 2023b)
LLaMAX2
PolyLM (Wei et al., 2023)
Yayi2 (Luo et al., 2023)
TowerInstruct (Alves et al., 2024)
Aya-23 (Aryabumi et al., 2024)
ChineseLLaMA2-Alpaca Cui et al. (2024)
LLaMA2-SFT (Taori et al., 2023)
Qwen2-7B-Instruct (Bai et al., 2023)
Swallow (Fujii et al., 2024)
B.2 Translation Models
@@
Appendix C The correlation between fertility and representation quality.
@@
Appendix D Introduction to KS-Lottery.
@@
Appendix E 1-hop translation in data augmentation is enough.
@@
Appendix F Design of parallel format
@@
Appendix G Comparison Results Between Our Model and GPT-4
@@
Appendix H Comparsion between LLaMAX2-Alpaca and language-specific LLMs.
In addition, we conduct a deeper analysis of translation output to identify the factors contributing to the limited improvement in translation performance. The experimental results in Table 12 indicate that the language-specific LLM obtained through the traditional pipeline tends to output specific languages, while LLaMAX2 can accurately produce the answer with the corresponding language.
@@
Appendix I Prompt Templates
We offer a comprehensive collection of prompt instruction templates, as illustrated in Table C, which are utilized for all evaluated LLMs. These templates are meticulously designed based on existing LLMs, playing a crucial role in obtaining accurate model results and ensuring fairness in comparisons. Our goal in providing these templates is to promote transparency and make it easier to reproduce our findings.